OpenStudy (anonymous):

The function g(x)=1x2-4x+4 is concave upwards on the intervals (-infinity, a) and (b, infinity) where a=? and b=?

OpenStudy (anonymous):

that should be 1/(x^2-4x+4), sorry!

OpenStudy (anonymous):

is this concavity problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\left(\begin{matrix}1\\ x^2-4x+4\end{matrix}\right) \] is this the correct format?

OpenStudy (turingtest):

finding concavity is a matter of evaluating the second derivative of a function, so g(x)=1/(x^2-4x+4)=(x^2-4x+4)^(-1) g''(x)=(-1)(2x-4)/(x-2)^2=-2/(x-2)=0 so we have a critical point at x=2 For x<2 like x=0 g''(0)=1>0 (concave up) For x>2 like x=4 g''(0)=-1/2<0 (concave down) so the interval (-infty,2) is concave down, while (2,infty) is concave up

OpenStudy (anonymous):

Oh okay. The fraction is what confused me.

OpenStudy (turingtest):

I actually made a slight mistake in the derivative, it should be: g(x)=1/(x^2-4x+4)=(x^2-4x+4)^(-1) g''(x)=(-1)(2x-4)/(x-2)^4=-2/(x-2)^3=0 so we have a critical point at x=2 For x<2 like x=0 g''(0)=1/8>0 (concave up) For x>2 like x=4 g''(4)=-1/8<0 (concave down) so the interval (-infty,2) is concave down, while (2,infty) is concave up so the answer doesn't change

OpenStudy (turingtest):

oh I'm sorry that's the first derivative, so we need to do the same thing for the next derivative. this should illustrate the process though.

OpenStudy (anonymous):

Thank you!!